Solve the initial value problem
y"+y=0 y(0)=2 y'(0)=0
1
Expert's answer
2020-12-08T19:07:54-0500

"y"+y=0,\,\, y(0)=2,\,\, y'(0)=0\\

\textsf{The auxiliary equation is}\\

m^2 + 1 = 0, m^2 = -1\\

m = \pm\sqrt{-1} = \pm j\\


\textsf{The general solution to the}\\ \textsf{differential equation is}\\

y = B\cos{x} + C\sin{x}\\


y(0) = B = 2, \therefore B = 2.\\


y' = -B\sin{x} + C\cos{x}\\


y'(0) = C = 0\\


\therefore \textsf{The particular solution is}\\


y = 2\cos{x}"